OpenStudy (anonymous):

A ball of mass 50 g tied to the end of a 50 cm inextensible string is whirled around in a vertical circle. Find the tension in the string when the ball is at the top of the circle

OpenStudy (yanasidlinskiy):

 T+mg=mv^2/r ------> Equation 1 r= length of the string (in m) m=mass of block in kg As I have stated the question V could be value of T